Presented here are summarized data from the balance sheets and income statements of Wiper, Inc.:

WIPER, INC.
Condensed Balance Sheets
December 31, 2017, 2016, 2015
(in millions)
2017 2016 2015
Current assets $ 707 $ 939 $ 793
Other assets 2,419 1,926 1,725
Total assets $ 3,126 $ 2,865 $ 2,518
Current liabilities $ 583 $ 836 $ 724
Long-term liabilities 1,530 997 870
Stockholders’ equity 1,013 1,032 924
Total liabilities and stockholders' equity $ 3,126 $ 2,865 $ 2,518
WIPER, INC
Selected Income Statement and Other Data
For the year Ended December 31, 2017 and 2016
(in millions)
2017 2016
Income statement data:
Sales $ 3,056 $ 2,919
Operating income 302 316
Interest expense 90 71
Net income 209 204
Other data:
Average number of common shares outstanding 41.9 47.3
Total dividends paid $ 56.0 $ 52.9

Required:

a. Calculate return on investment, based on net income and average total assets, for 2017 and 2016. (Do not round intermediate calculations. Round your answers to 1 decimal place.)

b. Calculate return on equity for 2017 and 2016. (Round your answers to 1 decimal place.)

c. Calculate working capital and the current ratio for each of the past three years. (Enter your answers in millions (i.e., 5,000,000 should be entered as 5). Round "Current ratio" to 1 decimal place.)

d. Calculate earnings per share for 2017 and 2016. (Round your answers to 2 decimal places.)

e. If Wiper's stock had a price/earnings ratio of 14 at the end of 2017, what was the market price of the stock? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

f. Calculate the cash dividend per share for 2017 and the dividend yield based on the market price calculated in part e. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

g. Calculate the dividend payout ratio for 2017. (Do not round intermediate calculations.)

h. Assume that accounts receivable at December 31, 2017, totaled $315 million. Calculate the number of days' sales in receivables at that date. (Use 365 days a year. Do not round intermediate calculations.)

i. Calculate Wiper's debt ratio and debt/equity ratio at December 31, 2017 and 2016. (Round "Debt ratio" to 1 decimal place and "Debt/equity ratio" to the nearest whole percent.)

j. Calculate the times interest earned ratio for 2017 and 2016. (Round your answers to 1 decimal place.)
Answer #1

Ans:

A) Return on investment(ROI)= net income/ average total assets *100

for 2017

average total assets= opening balance + closing balance /2

=$3126+$2865/2

=$5991/2

=2995.50

ROI= $209/2995.50*100

=0.0697*100

=6.97%

For 2016

average total assets= opening balance + closing balance /2

=$2865+$2518/2

=$5383/2

=2691.50

ROI= $209/2691.50*100

=0.0776*100

=7.76%

B Return on equity= Net income/ shareholder equity

For 2017= $209/1013

=0.2063*100

=20.63%

For 2016= $204/1032

=0.1976*100

= 19.76%

C working capital= Current Assets - Current liability

For 2017=$707 Million-$583 Million

=$124 Million or $124,000,000 or $124(as per question)

For 2016=$939-836

=$103 million or $103,000,000 or $103(as per question)

For 2015=$793-724

=$69 million or $ 69,000,000 or $69

Current ratio= current assets/ current liability

For 2017= $707/$583

=1.21:1

for 2016=$939/836

=1.12:1

For 2015=$793/724

=1.09:1

D Earnings per share=Net income- dividend paid/ weighted average share outstanding

For 2017=$209-$56/ 41.9

=$153/41.9

=$3.65

For 2016= $204-$52.9/47.3

=$151.1/47.3

= $3.19

G Dividend payout ratio= dividend paid/net income

for 2017=$56.0/209   

=0.26:1

H Number of days sale in account receivable=365/ account receivable ratio

account receivable= net credit sale/average inventory (given $315 at end of year so assume opening assume be 0)

=$3056/$315/2

=$3056/157.50

=19,40 times

number of days=365/19.40

=18.81 days

I Debt-equity ratio= debt /equity (for 2017)

=$1530/1013

=1.51:1

debt ratio=total liabilities/ total assets ( for 2017)

=$583+1530/$3126 (total liabilities= current liabilities+non current liabilities)

=$2113/$3126

=0.67:1

Debt-equity( for 2016) =$997/$1032

=0.966:1

debt ratio= $1833/$2865

=0.639:1

J times Interest earned ratio= income before interest and tax/ interest expense

for 2017=$302/90

=3.35:1

for 2016= $316/71

=4.45:1